# Beyond the Lens: Quantifying the Impact of Scientific Documentaries through Amazon Reviews

### v1

@@ 4.1. Classification Models
We leverage three types of classifiers for sentiment and impact classification: (1) three traditional machine learning algorithms, (2) two transformer-based models, and (3) five large language models. To create a train and test set, we split the data into 70% training and 30% test sets. For transformer-based models, we use 20% of the training data for validation. Stratified data split was used based on Impact labels to split the data due to an imbalance in the number of each group. We measured the classification performance on the test set after training each model using precision, recall, and F1-scores.
@@
We apply three baseline models, i.e., Support Vector Machine (SVM), Logistic Regression, and Decision Tree Classifier. We used Term Frequency-Inverse Document Frequency (TF-IDF) to vectorize the sentences.
@@
We leverage two transformer-based models, i.e., BERT (Devlin et al., 2018) and RoBERTa (Liu et al., 2019), to classify sentences. Models were fine-tuned using pre-trained weights from bert-base-uncased and roberta-base, respectively. The sentences were tokenized using the corresponding tokenizers (BertTokenizer or RobertaTokenizer). Each tokenized input was truncated or padded to a maximum sequence length of 512 tokens. Fine-tuning was performed using the Hugging Face API  [^4] https://huggingface.co/, with training arguments set to three epochs, a batch size of 8, a learning rate scheduler with 500 warmup steps, and a weight decay of 0.01. Model evaluation occurred at the end of each epoch using the validation set.
@@
We use GPT3.5 (OpenAI, 2023a) 555gpt-3.5-turbo-0125, GPT4 (OpenAI, 2023b) 666gpt-4-turbo-2024-04-09, GPT4o 777gpt-4o-2024-08-06, llama3*8b (AI@Meta, 2024) 888llama3-8b-8192 and Mixtral 8*7B (Jiang et al., 2024) 999mixtral-8x7b-32768 for classification in two different ways: with and without the full review as the textual context for the target sentence. In the first method (”w/o context”), the model is instructed to predict the correct label for the target sentence based on the definitions provided for each class. This is akin to having the model classify on only the top sentence (without the ”Full review”) depicted in Figure 1. In the second approach (with context), the full review is included in the prompt to guide the model in predicting the correct label for the target sentence. This is similar to providing the ”Full review” shown in Figure 1 and indicating the target sentence.
For these LLM-based classifiers, we do not use the training data. Instead, we use these models to predict the correct class for the test set using one of the mentioned prompting methods. This is possible due to the in-context learning abilities of advanced LLMs in understanding natural language and performing various linguistic tasks without the need for further training. Temperature is set to 0 for all LLM-models experiments for maximum consistency.
@@
To explore the helpfulness of additional fine-tuning for the classification purpose, we use the training set to fine-tune the models. Once the model is fine-tuned, we calculate the performance metrics on the test set again to assess whether fine-tuning has improved the classification performance.
@@ 5.1. Classification
Table 3 shows the performance of different classifiers on our dataset. Transformer- and LLM-based classifiers outperformed baseline models. GPT4 and GPT4o with context embedded in the prompts are performing the best compared to similar models, with F1 scores of 0.71 (GPT4o) for Impact and 0.85 (GPT4) for Sentiment, respectively. Our fine-tuned model of GPT4o resulted in the highest performance for both tasks, 0.76 and 0.88. This confirms the usefulness of fine-tuning LLMs with additional data. We also observe that while using full reviews as context is improving the performance of Sentiment classifiers, its helpfulness for Impact classification seems less significant.
@@ 5.2. Error Analysis
To enhance our understanding of our top-performing models, we conducted a detailed analysis of the fine-tuned GPT4o model based on the Impact class, which is the central focus of our study. Table 4 shows the F1 score per impact class. “Attitudes Toward the Film” has the highest classification performance of 0.86 while the F1 score for “Impersonal Report” is lowest at 0.42. This disparity may be linked to the imbalance of impact categories in the dataset (as seen across the entire dataset in Figure 2 and Table 2) during fine-tuning process or the more challenging nature of identifying this category.
Figure 3 demonstrates the normalized confusion matrix of Impact classes, where we can see the most common type of misclassifications produced by the model. It is evident that “Attitudes Toward the Film” and “Impersonal Report” are most often misclassified as “Shift in Cognition”, while “Interest with Science Topic”, “Not applicable”, and “Shift in Cognition” are mostly misclassified as “Attitudes Toward the Film”.
@@ 5.3. Generalizability Test
To evaluate the generalizability of our models in successfully capturing the impact of movies, we apply the best model, defined as having the highest classification performance (”GPT4o Fine-tuned” in Table 3) to the Hubble dataset and predict the labels for both Impact and Sentiment.
Table 6 shows the distribution of predicted classes for this dataset. While the majority of reviews show “Positive” sentiment toward the Hubble documentary, “Attitudes Toward the Film” is the most frequent type of impact predicted by our model in the movie review. This aligns with the prevalence of the “Attitudes Toward the Film” impact category among the other movies in Figure 1 and Table 2.
[…]
@@ 5.4. Thematic Analysis
Table 5 shows the themes of discussions in sentences, labeled with each impact category using LLooM. We exclude “Not applicable” from Table 5 as this category covers a wide variety of comments which are not relevant to the overall analysis of the movies’ impacts.
Each Impact contains several related themes based on how reviewers have described the movie. While sentences labeled as “Attitudes Toward the Film” discuss various positive, negative, or mixed feedback on the movie, reviews tagged with “Shift in Cognition” cover themes such as space science, weather, and environmental concerns or information quality. Context around “Interest with Science Topic” includes mentions of science and technical aspects, emotional and experiential reactions, content quality, and learning; while “Impersonal Report” often consists of safety issues or historical and informational aspects of the movies. These themes collectively show how the viewers have interacted with scientific documentaries and how these movies engaged with their cognition, attitudes, and interests.
Figure 3. Confusion Matrix of predicted labels for Impact classes using GPT4o finetuned model
Attitudes Toward the Film 0.86
Impersonal Report 0.42
Interest with Science Topic 0.52
Not applicable 0.65
Shift in Cognition 0.62
Table 4. F1 score per class for Impact
@@
Impact Sentiment
Attitudes Toward the Film 307 Positive 250
Not applicable 50 Neutral 94
Interest with Science Topic 17 Negative 56
Shift in Cognition 15
Impersonal Report 11
Table 6. Distribution of predicted classes for Hubble dataset
@@ 6. Discussion
In conversation with the CSVT  several of the interests the group emphasized were (1) a desire to see a breakdown between positive and negative views of their documentaries, (2) an understanding of what film attributes might drive a viewer’s opinion of the film, and (3) that, as often engagement in and of itself is a goal, what measurements of audience engagement might be evident in comments. These interests align with our research questions (“RQ1” and “RQ2”).
RQ1: The most prevalent impacts of scientific documentaries.
Our findings highlight the substantial role that scientific documentaries play in shaping audience engagement and attitudes. The predominance of positive sentiments, particularly those linked to “Attitudes Toward the Film” (37% of all sentences), suggests that these films successfully resonate with their audiences, likely due to the compelling storytelling and production quality typical of impactful documentaries. This supports existing literature on the ability of documentary films to evoke emotional responses and create a sense of connection between viewers and experts (Nisbet and Aufderheide, 2009; Corner, 2002).
[…]
RQ2: Common themes and discourse in viewers’ feedback.
As indicated in Table 5, there are several themes that align with each impact category. For example, attitudes toward the film which lean negative often mention poor narration or production quality or a lack of excitement, while positive attitudes mention a desire for similar content and that the content is engaging and informative. These thematic trends align with interest (2) of the CSVT.
Additionally, many themes listed in Table 5 involve engagement as well as the visualization/technical aspects of the films, aligning with interest (3) of the CSVT. Not only is engagement talk more evidence of learning, but it is also a building block that can lead to behaviors with prosocial impact, like de-stigmatizing and sharing science information with community in informal environments, improving personal science identity which can lead to the growth of future science professionals, and advocating for more research funding, political support, and science-motivated decision making (e.g., Lee-Robbins and Adar, 2022b).
@@ 7. Conclusion and Future Work
Beyond the immediate goals of the CSVT the results of our best model as outlined in Table 3 and Section 5.1 indicate it is likely the model could be utilized on future productions by the CSVT and other teams in order to estimate the impact of their produced works. Additionally, the results of our generalizability analysis presented in Table 6 and Section 5.3 indicate the model should be applicable to a wide range of scientific CSV-style documentaries.
While our model was trained and tested on Amazon comments, our collaboration also has access to YouTube comments for a subset of the analyzed films. Future work will include an analysis of these comments – there is evidence of a bias toward more negative reviews on YouTube (e.g. Ardestani, 2024), making such a comparison a test of the inter-platform generalizability of our model.
The CSVT expressed interest in tying specific learning goals to sentiment and impact categories and/or impact themes. As learning goals are often developed to guide the creation of these scientific documentaries and are integral to grant funding, our large-scale quantitative analysis of the impact of these scientific documentaries would be a useful complement to the smaller-scale audience-based feedback the CSVT currently collects. While beyond the scope of the present work, future research with a larger dataset of cross-platform and multilingual input could benefit from this more detailed analysis.
@@
Overall, our results underscore the multifaceted impact of scientific documentaries. They engage audiences emotionally, provoke intellectual curiosity, and influence perceptions, making them a powerful medium for science communication. Our findings contribute to a growing body of evidence on the effectiveness of media in bridging the gap between science and the public, emphasizing the value of leveraging documentaries to achieve broader educational and societal goals. Future research could explore the long-term impacts of such engagement, particularly in terms of sustained interest in science and changes in behavior or policy advocacy.

### v2

@@ 4.1. Classification Models
We leveraged three types of classifiers for sentiment and impact classification: (1) three traditional machine learning algorithms, (2) two transformer-based models, and (3) five large language models. To create a train and test set, we split the data into 70% training and 30% test sets. For transformer-based models, we use 20% of the training data for validation. Stratified data split was used based on Impact labels to split the data. We measured the classification performance on the test set after training each model using precision, recall, and F1 scores.
@@
We applied three baseline models, i.e., Support Vector Machine (SVM), Logistic Regression, and Decision Tree Classifier. We used Term Frequency-Inverse Document Frequency (TF-IDF) to vectorize the sentences.
@@
We leveraged two transformer-based models, i.e., BERT (Devlin et al., 2018) and RoBERTa (Liu et al., 2019), to classify sentences. Models were fine-tuned using pre-trained weights from bert-base-uncased and roberta-base, respectively. The sentences were tokenized using the corresponding tokenizers (BertTokenizer or RobertaTokenizer). Fine-tuning was performed using the Hugging Face API [^4] https://huggingface.co/, and hyperparameters were set to three epochs, a batch size of 8, a learning rate scheduler with 500 warmup steps, and a weight decay of 0.01. Model evaluation occurred at the end of each epoch using the validation set.
@@
We used three closed-weight models from OpenAI: GPT3.5 (gpt-3.5-turbo-0125) (OpenAI, 2023a), GPT4 (gpt-4-turbo-2024-04-09) (OpenAI, 2023b), GPT4o (gpt-4o-2024-08-06), and two open-weight LLMs: llama3*8b (llama3-8b-8192) (AI@Meta, 2024) and Mixtral 8*7B (mixtral-8x7b-32768) (Jiang et al., 2024) for classification in two different ways: with and without the context (full review) as the textual context for the target sentence. In the first method (“w/o context”), the model is instructed to predict the correct label for the target sentence based on the definitions provided for each class. This is akin to having the model classify on only the top sentence (without the “Full review”) depicted in Figure 2. In the second approach (“with context”), the full review is included in the prompt to guide the model in predicting the correct label for the target sentence. This is similar to providing the “Full review” shown in Figure 2 and indicating the target sentence. For the LLM-based classifiers, we used the test set to evaluate the models’ performance as well as the efficiency of each prompting method. Temperature was set to 0 for all model experiments for maximum consistency.
@@
To explore the helpfulness of additional fine-tuning for the classification purpose, we finetuned GPT4o (gpt-4o-2024-08-06) using the training set. After fine-tuning, we repeated the classification task with the same prompts to predict the label for each sentence. The performance was then evaluated on the test set to assess the improvement compared to the base model.
@@ 5.1. Classification
Table 3 shows the performance of different classifiers on our dataset. Transformer- and LLM-based classifiers outperformed baseline models. GPT4 and GPT4o with context embedded in the prompts are performing the best compared to similar models, with F1 scores of 0.71 (GPT4o) for Impact and 0.85 (GPT4) for Sentiment, respectively. Our fine-tuned model of GPT4o resulted in the highest performance for both tasks, 0.76 and 0.88. This confirms the usefulness of fine-tuning LLMs with additional data. We also observed that while using full reviews as context is improving the performance of Sentiment classifiers, its helpfulness for Impact classification seems less significant.
@@ 5.2. Error Analysis
To enhance our understanding of our top-performing models, we conducted a detailed analysis of the fine-tuned GPT4o model based on the Impact class, which is the central focus of our study. Table 5 shows the F1 score per impact class (center column). “Attitudes Toward the Film” has the highest classification performance of 0.86 while the F1 score for “Impersonal Report” is lowest at 0.42. These trends rightly align with the Cohen’s Kappa measure of human annotator intercoder agreement (McHugh, 2012) shown in the right column of Table 5. While most categories show substantial agreement between annotators ($\kappa>0.6$), “Impersonal Report” is on the borderline between moderate and substantial agreement. The relatively larger disparity in F1 score from the model’s classifications when compared to differences in Cohen’s Kappa may be linked to the imbalance of impact categories in the dataset (as seen across the entire dataset in Figure 1 and Table 2) during fine-tuning process or the more challenging nature of identifying this category.
Table 4 demonstrates the confusion matrix of Impact classes, where we can see the most common type of misclassifications by the model. It is evident that “Attitudes Toward the Film” and “Impersonal Report” are most often misclassified as “Shift in Cognition”, while “Interest with Science Topic”, “Not applicable”, and “Shift in Cognition” are mostly misclassified as “Attitudes Toward the Film”.
True Label Predicted Label
Attitudes Impersonal Interest NA Shift
Attitudes Toward the Film 202 1 4 5 17
Impersonal Report 3 4 0 1 4
Interest with Science Topic 10 0 11 1 4
Not applicable 20 2 0 31 4
Shift in Cognition 8 0 1 1 32
Table 4. Confusion Matrix for True and Predicted Impact Categories: Attitudes (Attitudes Toward the Film), Impersonal (Impersonal Report), Interest (Interest with Science Topic), NA (Not applicable), and Shift (Shift in Cognition)
Impact Category F1 Cohen’s Kappa
Attitudes Toward the Film 0.86 0.68
Impersonal Report 0.42 0.61
Interest with Science Topic 0.52 0.67
Shift in Cognition 0.62 0.67
Not applicable 0.65 0.67
Table 5. F1 score (model) and Cohen’s Kappa (human-annotations) per class for Impact.
@@ 5.3. Generalizability Test
To evaluate the generalizability of our models in successfully capturing the impact of movies, we applied the best model, with the highest classification performance (“GPT4o Fine-tuned”) to the Hubble dataset and predicted the labels for both Impact and Sentiment. Table 6 shows the distribution of predicted classes for this dataset. While the majority of reviews show “Positive” sentiment toward the Hubble documentary, “Attitudes Toward the Film” is the most frequent type of impact predicted by our model in the movie review. This aligns with the prevalence of the “Attitudes Toward the Film” impact category among the other movies in Table 2.
To assess the model’s performance, we tasked two annotators with evaluating the accuracy of the predicted labels for both the Impact and Sentiment categories. The mean agreement for Impact is approximately 85.3%. For Sentiment agreement is $\sim$87.6%, demonstrating a substantial level of alignment between the annotators and the classification model.
Impact Sentiment
Attitudes Toward the Film 307 Positive 250
Not applicable 50 Neutral 94
Interest with Science Topic 17 Negative 56
Shift in Cognition 15
Impersonal Report 11
Table 6. Distribution of predicted classes for Hubble dataset
@@ 5.4. Thematic Analysis
Table 7 shows the themes of discussions in sentences, labeled with each impact category using LLooM. We exclude “Not applicable” from Table 7 as this category covers a wide variety of comments which are not relevant to the overall analysis of the movies’ impacts. Each Impact contains several related themes based on how reviewers have described the movie. While sentences labeled as “Attitudes Toward the Film” discuss various positive, negative, or mixed feedback on the movie, e.g., “Kind of boring bearly any action” or “The graphics are… OUT OF THIS WORLD”, reviews tagged with “Shift in Cognition” cover themes such as space science and environmental concerns or information quality, e.g., “Jupiter is a stunningly gorgeous planet, and it’s amazing to finally learn some of its secrets.”
Context around “Interest with Science Topic” includes mentions of science and technical aspects, emotional and experiential reactions, content quality, and learning, e.g., “If observing and learning of the universe is your thing, you will love this”. “Impersonal Report” often consists of safety issues or historical and informational aspects of the movies, e.g., “An 1859 type event will occur again and cause 1 to 2 Trillion dollars in damage.” These themes collectively show how the viewers have interacted with scientific documentaries and how these movies engage with their cognition, attitudes, and interests.
@@ 6. Discussion
In conversation with the AVL  several of the interests the group emphasized were (1) a desire to see a breakdown between positive and negative views of their documentaries, (2) an understanding of what film attributes might drive a viewer’s opinion of the film, and (3) that, as often engagement in and of itself is a goal, what measurements of audience engagement might be evident in comments. These interests align with our two research questions.
RQ1: The most prevalent impacts of scientific documentaries. Our findings highlight the substantial role that scientific documentaries play in shaping audience engagement and attitudes. The predominance of positive sentiments, particularly those linked to “Attitudes Toward the Film” (37% of all sentences), suggests that these films successfully resonate with their audiences. This is likely due to the compelling storytelling and production quality typical of impactful documentaries. This supports existing literature on the ability of documentary films to evoke emotional responses and create a sense of connection between viewers and experts (Nisbet and Aufderheide, 2009; Corner, 2002).
The disparity between “Positive”/“Attitudes Toward the Film” (37%) and “Negative”/“Attitudes Toward the Film” (21.9%) further underscores a general trend of audience favorability. This aligns with findings that highlight how carefully crafted documentaries can frame scientific content in ways that promote optimism, curiosity, and a sense of wonder about the natural or scientific world (Gregory and Miller, 1998). The approximately $\sim$1.7 greater prevalence of positive over negative sentiments reflects the potential of documentaries to act as persuasive and motivational tools, aligning with interests (1) and (3) of the AVL, focusing on emotional resonance and knowledge dissemination. “Interest with Science Topic” (5.2%) and “Shift in Cognition” (6.9%) prevalence, as significant secondary impacts, indicate that these films go beyond entertainment, fostering intellectual engagement and engaging with viewer perspectives. This mirrors findings from studies on science communication that emphasize the dual purpose of documentaries: sparking curiosity and promoting critical thinking (Dahlstrom, 2014). The fact that these impact types, although not the most prevalent, still contribute to a notable portion of the dataset shows that documentaries can serve as informal educational tools, stimulating deeper engagement with scientific topics and potentially influencing public understanding of science.
RQ2: Common themes and discourse in viewers’ feedback. As indicated in Table 7, there are several themes that align with each impact category. For example, attitudes toward the film which lean negative often mention poor narration or production quality or a lack of excitement, while positive attitudes mention a desire for similar content and that the content is engaging and informative. These thematic trends align with interest (2) of the AVL.
Additionally, many themes listed in Table 7 involve engagement as well as the visualization/technical aspects of the films, aligning with the interest (3) of the AVL. Not only is engagement talk more evidence of learning, but it is also a building block that can lead to behaviors with prosocial impact. These behaviors include de-stigmatizing and sharing science information with the community in informal environments, improving personal science identity (which can lead to the growth of future science professionals), and advocating for more research funding, political support, and science-motivated decision-making (e.g., Lee-Robbins and Adar, 2022b).
@@ 7. Conclusion and Future Work
This paper analyzes the impact of scientific films using user-generated reviews and measures the impact using a novel taxonomy of change and engagement. Beyond the immediate goals of AVL the results of our best model as outlined in Table 3 and Section 5.1 indicate our model’s efficiency in predicting impact and sentiment which can be useful in future productions assessment by the AVL and other teams to assess the impact of their produced works. Additionally, the results of our generalizability analysis presented in Table 6 and Section 5.3 indicate the model should be applicable to a wide range of scientific CSV-style documentaries. Overall, our results underscore the multifaceted impact of scientific documentaries. They engage audiences emotionally, provoke intellectual curiosity, and influence perceptions, making them a powerful medium for science communication. Our findings contribute to a growing body of evidence on the effectiveness of media in bridging the gap between science and the public, emphasizing the value of leveraging documentaries to achieve broader educational and societal goals. Future research could explore the long-term impacts of such engagement, particularly in terms of sustained interest in science and changes in behavior or policy advocacy.
The AVL expressed interest in tying specific learning goals to sentiment and impact categories and/or impact themes. As learning goals are often developed to guide the creation of these scientific documentaries and are integral to grant funding, our large-scale quantitative analysis of the impact of these documentaries would be a useful complement to the smaller-scale audience-based feedback the AVL currently collects. While our model was trained and tested on Amazon comments, our collaboration also has access to YouTube comments for a subset of the analyzed films. Future work will include an analysis of these comments, however, there is evidence of a bias toward more negative reviews on YouTube (e.g. Ardestani, 2024), making such a comparison a test of the inter-platform generalizability of our model.
8. Limitations and Ethics Statement
Our data collection is limited only to English reviews, impacting the generalizability of our findings. We also limited the scope of our study to only one platform, i.e., Amazon. This can enable platform-specific biases to influence the findings, and classification performance and cause potential inaccuracies. In addition, media impact reviews are self-reported and collected only once. Without access to follow-up reviews or more nuanced, contextualized reports, we cannot be certain of the persistence and durability of such reported impacts. Finally, we acknowledge that LLM-based classifiers often make mistakes in predicting the best label, resulting in misclassified reviews. Future improvements in these models and better prompt engineering can enhance the performance of LLM classifiers.
Acknowledgements.
The authors thank the work of Alistair Nunn, Rishabh Sharma, and Shashwat Mann for their annotation and analysis consultations. The authors additionally thank the members of the Advanced Visualization Laboratory at the National Center for Supercomputing Applications for their collaboration on this work. Finally, the authors thank the referees of this work for their valuable input.
Appendix A Annotation Platform
We used the Zooniverse for the annotation process. Zooniverse boasts over one million active members collaborating on hundreds of projects ranging from the space sciences to transcription of historical documents [^5] https://www.zooniverse.org/about. Here, “first choice” is defined as the strongest sentiment expressed in the statement. If there were two equally strong sentiments expressed, then the sentiment that appeared first was selected as the “first choice” (the analysis of “second-choice” selections is relegated to a future paper). Once the user made their selection, they could select “Done” to move to the next stage of the annotation (green button in Figure 2) or “Done & Talk” to ask questions in the main forum (blue button in Figure 2). The annotator was then prompted to repeat the process with their first choice for impact category (categories listed in the first column of Table 1).
Figure 2. First stage of annotation process in which users are instructed to select their first choice for the sentiment category for each sentence. Each sentence (top) is shown in the context of the full review (bottom). The annotator moves to the next stage of annotation with the green “Done” button (or blue “Done & Talk” button to post comments in the project forum). More information (comment URL and film title) is accessed with the circled “i” Info button and annotation can be saved in the user’s personal collection for later reference with the heart Favorite button.
Appendix B Prompt design
Figure 3. Prompt used for prediction of impact categories
Figure 3 illustrates the prompt we used to develop LLM classifiers for impact categories that are instructed to label a target sentence, using full reviews as context. The prompt begins with an instruction of the classification task, followed by the provided definitions of impact labels, and feeding both the context and target sentence to the prompt at the end.